OpenStudy (anonymous):

What is square root of 114 approximated to the nearest tenth?

OpenStudy (anonymous):

I'll medal!

OpenStudy (freckles):

sounds like you should get your calculator out and put in sqrt(114).

OpenStudy (freckles):

the hundredths digit will tell you to round or leave the tenth digit the same

OpenStudy (anonymous):

okay, thanks

OpenStudy (freckles):

Do you want to tell me what you get after using the calculator and rounding to the nearest tenth?

OpenStudy (anonymous):

sure

OpenStudy (freckles):

I can also show you how to approximate sqrt(114) without a calculator if you are interested

OpenStudy (anonymous):

That would be cool, I suck at math

OpenStudy (freckles):

ok so you let's think about just the number inside the sqrt the 114 that is Let's think what two nearest perfect squares is 114 in between. 100 is a perfect square since 10*10=100 121 is a perfect square since 11*11=121 114 is certainly in between 100 and 121 that is we have the following inequality \[100 <114<121\]

OpenStudy (freckles):

now the difference between 121 and 100 is 21 and the difference between 114 and 100 is 14 so 114 is 14/21 the distance between 100 and 121 ok let's keep that in mind \[100 <114<121\] take square root \[10<\sqrt{114} <11\] so we know that \[\sqrt{114} \text{ is approximately } 10+\frac{14}{21}\]

OpenStudy (freckles):

you know because we already knew it was 10 something since 10.something is between 10 and 11

OpenStudy (freckles):

so you can write 14/21 as a decimal then add the 10 then round to the nearest 10ths

OpenStudy (freckles):

that is an algebraic approach to approximate square roots there are also other ways to approximate it like using calculus

OpenStudy (anonymous):

thank you!

OpenStudy (freckles):

np